OpenStudy (anonymous):

6(4z+3)-10(7z-1)=3(2-19z) Solve for z

OpenStudy (mathmale):

You have shared only one equation here, and that equation does not involve x at all. How, then, can you expect to solve for x and y? Is it possible that there was another equation that should have come with this one?

OpenStudy (anonymous):

Oh no,I had post a wrong equation,may I edit

OpenStudy (mathmale):

type everything you have all over again, unless you know how to edit your previous post.

OpenStudy (anonymous):

But I already edit before your replied,sorry I edit before your replied.

OpenStudy (mathmale):

6(4z+3)-10(7z-1)=3(2-19z). Multiply out each product. For example, 6(4z+3)=24z+18.

OpenStudy (mathmale):

then group all the z terms on the left side and all the constants on the other side.

OpenStudy (anonymous):

Is z=-2?

OpenStudy (mathmale):

I need to see your actual work, if you want me to check your work. Multiply out each product. For example, 6(4z+3)=24z+18.

OpenStudy (anonymous):

6(4z+3)-10(7z-1)=3(2-19z) =24z+18-70z+10=6-57z =24z -70z+18+10=6-57z =-46+28=6-57z 11z+28=6 11z=-22 z=-2

OpenStudy (anonymous):

sorry I mean -46z

OpenStudy (anonymous):

24z+18-70z+10=6-57z -46z+28=6-57z -46z+57z=6-28 11z=-22 Z=-2